# WeQA: A Benchmark for Retrieval Augmented Generation in Wind Energy Domain

### v2

@@ Abstract
In the rapidly evolving landscape of Natural Language Processing (NLP) and text generation, the emergence of Retrieval Augmented Generation (RAG) presents a promising avenue for improving the quality and reliability of generated text by leveraging information retrieved from user specified database. Benchmarking is essential to evaluate and compare the performance of the different RAG configurations in terms of retriever and generator, providing insights into their effectiveness, scalability, and suitability for the specific domain and applications. In this paper, we present a comprehensive framework to generate a domain relevant RAG benchmark. Our framework is based on automatic question-answer generation with Human (domain experts)-AI (Large Language Model (LLM)) teaming. As a case study, we demonstrate the framework by introducing WeQA, a first-of-its-kind benchmark on the wind energy domain which comprises of multiple scientific documents/reports related to environmental impact of wind energy projects. Our framework systematically evaluates RAG performance using diverse metrics and multiple question types with varying complexity level. We also demonstrate the performance of different models on our benchmark.
@@ 1 Introduction
In recent years, the advancements in LLM have revolutionized various natural language processing tasks, including text and response generation. However, text generation using LLM often encounters challenges such as generating irrelevant or incoherent outputs, perpetuating biases ingrained in the training data, and struggling to maintain context and factual accuracy. These issues pose significant obstacles to achieving human-level performance in automated text generation systems. RAG effectively mitigates these common challenges by incorporating retrieved information to enhance coherence and factual accuracy, thus minimizing the generation of fictitious or irrelevant content Gao et al. (2024); Lewis et al. (2021). Furthermore, concurrent works suggest RAG is the most sought approach for adapting models towards niche scientific domain such as nuclear, renewable energy, environmental policy, etc. Munikoti et al. (2024a, b); Phan et al. (2023)
As this innovative approach gains traction within the research community and industry applications, its effectiveness and robustness must be rigorously assessed against established benchmarks to ensure its practical utility and reliability Chen et al. (2023a). Benchmarking is essential to establish standardized evaluation metrics and dataset that can effectively capture the nuances of text quality, coherence, factual accuracy, and relevance. Further, it facilitates comparison between RAG and existing text generation methods, shedding light on its strengths, limitations, and potential areas for improvement Xiong et al. (2024). A robust benchmarking framework can enable researchers and practitioners to systematically investigate the impact of various parameters, such as retrieval strategies, model architectures, and training data, on the performance of RAG Ray (2023).
@@
In benchmarking RAG for text generation, it is crucial to evaluate its performance across a diverse set of questions to ensure its efficacy in handling various linguistic contexts and user intents Lyu et al. (2024). A set of well curated and diverse questions enable a comprehensive assessment of RAG’s ability to generate coherent and relevant responses across various domains, ensuring its applicability in real-world scenarios. To generate such questions, automated methods leveraging NLP techniques can be employed. These methods include rule-based approaches, template filling, and neural network-based models, which enable the efficient creation of diverse question sets by leveraging linguistic patterns and semantic transformations.
@@
Human-curated questions offer a level of linguistic richness and contextual relevance that may be challenging to achieve solely through automated generation methods Zhang et al. (2024). By leveraging human expertise and domain knowledge, curated question sets can encompass a broader spectrum of linguistic variations, domain-specific considerations, and nuanced semantics, providing a more comprehensive evaluation of RAG’s performance across diverse scenarios and applications. Combining automated generation with human curation for benchmarking RAG offers a synergistic approach to ensure both efficiency and quality in question sets. This hybrid approach leverages the strengths of both automated and human-driven processes, that provide efficient and robust evaluation metrics for RAG’s performance.
@@
In this work, we present a hybrid workflow to benchmark RAG s, which combines rapid question generation through automated methods, augmented with properly designed human prompts to generate diverse set of questions. Our proposed benchmarking framework is used to generate questions from documents related to wind energy domain. These questions serve as a tool to evaluate the performance of RAG-based LLM s, which are designed to answer queries related to these extensive and comprehensive documents. Given the vast amount of information contained in these documents, manually reviewing them is impractical, making RAG-based LLM s essential for generating accurate responses to specific queries. Our benchmarking framework assesses the effectiveness of these models in accurately retrieving and responding to queries, ensuring that they can reliably process and provide relevant information from the documents.
Contributions  The paper introduces a novel benchmark in a specific domain and also proposes a generic framework to evaluate the performance of RAG-based LLM s in responding to different types of questions. This framework is designed to be adaptable across various domains, with a specific focus on wind energy-related documents in this study. The contributions of this research are as follows: (i) We present WeQA,  [^1] This benchmark will be made publicly available. the first benchmark in the Wind Energy domain, (ii) our proposed framework is domain-agnostic, so it can be tailored for any desired niche domain (iii) we introduce a hybrid method to automatically generate various types of questions, producing both objective and subjective responses. The framework also generates questions from different sections of documents to evaluate LLM performance across various sections and question types, and (iv) we utilize existing scoring frameworks like $\mathsf{RAGAS}$ to evaluate model performance, incorporating different LLM s as evaluators for scoring. This approach ensures scalability and quick reproducibility of this approach, while also providing a holistic comparison of LLM performance in terms of responding to questions and assessing or comparing LLM responses with the ground truth answers.
Figure 1: An overview of data extraction and curation pipeline to generate a vector database of relevant wind energy related documents.
@@ 2 Related Works
The Stanford Question Answering Dataset (SQuAD) Rajpurkar et al. (2016), consisting of 100,000+ questions and a reading comprehension dataset, is arguably the most famous general QA benchmark of the field. They contain three subtasks within QA: reading comprehension, Open-domain QA, and missing word prediction. The AI2 Reasoning Challenge (ARC) Clark et al. (2018) is another major work, which contains almost $8,000$ science questions in English, and also included questions that neither a retrieval-based algorithm nor a word co-occurrence algorithm were able to answer correctly. Similarly, the MCTest dataset Richardson et al. (2013) consists of $500$ stories and $2000$ young children level multiple-choice reading comprehension questions. Some other notable QA benchmarks include Big Bench Srivastava et al. (2022), ARC2 Bhakthavatsalam et al. (2021), GLUE Wang et al. (2018), CommonsenseQA Talmor et al. (2018), TriviaQA: 650K QA pairs with evidence Joshi et al. (2017), Search QA Dunn et al. (2017), NewsQA: 10K news articles Trischler et al. (2016), inter alia.
More recently, there have been several benchmarks that focus on scientific and adjacent fields. The scientific portions of the MMLU Hendrycks et al. (2020) benchmark is perhaps one of the most widely used science benchmarks, which include college and high school level questions in Physics, Chemistry, Biology, Computer Science, etc. Science Questions: 1K multiple choice questions in AI2R Talmor et al. (2018) and SciQ Dataset: Welbl et al. (2017) 13,679 multiple choice science questions are two other major benchmarks in the scientific domain, as is the SciQA Auer et al. (2023), a scientific QA benchmark created by using knowledge graphs of academic articles. SciRepEvalSingh et al. (2022) is a benchmark that has four different task types – classification, regression, proximity – over scientific documents.
Similarly, some of the other most recent works include SciBenchWang et al. (2023), a benchmark of $\sim$700 questions sourced from textbooks for college-level science problems and QASA Lee et al. (2023), a QA benchmark of $\sim$1800 questions to test reasoning on scientific articles, specifically in AI and ML domains. There are also benchmarks that address specific fields, with TheoremQA Chen et al. (2023b) for mathematics, emrQA Pampari et al. (2018) for medicine, and BioRead Pappas et al. (2018) and BioMRC Pappas et al. (2020) for biology, and NukeBERT Jain et al. (2020) and NuclearQA Acharya et al. (2023) for the nuclear domain.
While these scientific domains are related to our task in terms of technological similarity, to our knowledge, there are no benchmarks for our specific field and this is the first such work. The only one that comes close is the NEPAQuAD benchmark Phan et al. (2023) that deals with QA task for Environmental Impact Statement (EIS) documents.
@@ 3 Dataset Creation
We constructed a data extraction and curation pipeline to extract text, image, and table information from wind energy-related documents as depicted in Figure 1. Utilizing large language model (LLM) based methods such as the Unstructured.io tool Raymond (2023), we efficiently extracted information and converted it into JSON elements. These JSON elements were then organized into a schema, creating a page-wise assortment of text, table, and image elements. This structured format ensures that the extracted data is easily accessible and can be accurately referenced during model training and evaluation.
@@ 4 Methodology
Next, we present two approaches for the hybrid question generation procedure required for LLM benchmarking purposes. The first approach engineers the prompt to generate well curated enhanced quality questions. The second approach summarizes the provided text chunks and generates questions from the summaries.
Hybrid prompts.  We use GPT-4 to automatically generate questions from a given text chunk by providing particular prompts for each question type. The prompt is structured as follows:
Generate {num} questions given the content provided in the following paragraph. Restrict the type of questions to {question type} questions.
{Text chunk to generate the questions.}
@@
Figure 3: Summary of “introduction” section of a report Invenergy (2014) generated by GPT-4. The question and the answer are generated from the summarized text chunk. The answer is retrieved from a subset of text in the chunk, shown here in red.
An important aspect of our approach is curating the automatically generated questions to enhance the quality. To this end, we manually identify questions which are best suited for the purpose of benchmarking LLM s. We perform this process for each type of question, so that we include particular grammatical structures for each question type. Thereafter, we use these identified questions as few-shot examples to regenerate questions using the automatic question generation framework. The updated prompt looks as follows:

### v3

@@ Abstract
Wind energy project assessments present significant challenges for decision-makers, who must navigate and synthesize hundreds of pages of environmental and scientific documentation. These documents often span different regions and project scales, covering multiple domains of expertise. This process traditionally demands immense time and specialized knowledge from decision-makers. The advent of Large Language Model (LLM) s and Retrieval Augmented Generation (RAG) approaches offer a transformative solution, enabling rapid, accurate cross-document information retrieval and synthesis. As the landscape of Natural Language Processing (NLP) and text generation continues to evolve, benchmarking becomes essential to evaluate and compare the performance of different RAG-based LLM s. In this paper, we present a comprehensive framework to generate a domain relevant RAG benchmark. Our framework is based on automatic question-answer generation with Human (domain experts)-AI (LLM) teaming. As a case study, we demonstrate the framework by introducing WeQA, a first-of-its-kind benchmark on the wind energy domain which comprises of multiple scientific documents/reports related to environmental aspects of wind energy projects. Our framework systematically evaluates RAG performance using diverse metrics and multiple question types with varying complexity level, providing a foundation for rigorous assessment of RAG-based systems in complex scientific domains and enabling researchers to identify areas for improvement in domain-specific applications.
@@ 1 Introduction
In recent years, the advancements in LLM have revolutionized various natural language processing tasks, including text and response generation. However, text generation using LLM often encounters challenges such as generating irrelevant or incoherent outputs, perpetuating biases ingrained in the training data, and struggling to maintain context and factual accuracy Wu et al. (2024). These issues pose significant obstacles to achieving human-level performance in automated text generation systems. RAG effectively mitigates these common challenges by incorporating retrieved information to enhance coherence and factual accuracy, thus minimizing the generation of fictitious or irrelevant content Gao et al. (2024); Lewis et al. (2021). Furthermore, concurrent works suggest RAG is the most sought approach for adapting models towards accelerating repetitive and data intensive tasks in niche scientific domain such as nuclear, renewable energy, environmental policy, etc. Munikoti et al. (2024a, b); Phan et al. (2023). While RAG-based systems have demonstrated promising capabilities in streamlining document analysis tasks across various professional domains, their integration into critical decision-making processes like permitting wind energy projects remains constrained due to legitimate concerns about trust and reliability.
In this work, we create benchmarks to assess RAG-based LLM performance in the domain of permitting wind energy projects. Environmental Impact Statements (EIS) represent the cornerstone documentation within this permitting landscape, serving as comprehensive analyses that evaluate the potential environmental consequences of proposed wind energy developments. These documents play a pivotal role in promoting informed decision-making by ensuring transparency and incorporating diverse stakeholder perspectives into the approval process Bond et al. (2024). By providing detailed evaluations of environmental effects, alternatives analysis, and mitigation measures, EIS documentation facilitates the responsible development of wind energy infrastructure while building public trust at the same time.
As RAG-based LLM s gain traction for domain-specific applications such as wind energy permitting, their effectiveness must be rigorously assessed through robust benchmarks to ensure its practical utility and reliability Chen et al. (2023a). Establishing high-quality benchmarks is essential to evaluate their abilities to perform regulatory-focused reasoning, accurately interpret complex EIS documents, and support logical deductions grounded in the documents. Such benchmarks facilitate systematic assessment of how well RAG-based LLM s can handle the nuanced requirements of the domain Xiong et al. (2024). A robust evaluation framework allows researchers and practitioners to investigate the impact of retrieval strategies, model architectures, and training data, on the performance of RAG, while building confidence in automated tools for critical environmental decision making Ray (2023).
@@
In benchmarking RAG for wind energy project permitting applications, it is crucial to evaluate its performance across a diverse set of questions that reflect the complexity and variability of real-world permitting scenarios Lyu et al. (2024). A set of well curated and diverse questions enable a comprehensive assessment of RAG’s ability to interpret EIS documents, analyze environmental impacts, evaluate regulatory compliance, and generate coherent responses to permitting-related queries that practitioners encounter during wind energy project review processes. To generate such questions, automated methods leveraging NLP techniques can be employed, including rule-based approaches that capture language patterns from relevant documents, template filling methods that incorporate wind energy terminologies, and neural network-based models that can efficiently create diverse question sets by leveraging the semantic relationships inherent in EIS and other documents related to wind energy projects.
@@
Human-curated questions offer a level of linguistic richness and contextual relevance that may be challenging to achieve solely through automated generation methods, particularly in specialized domains such as wind energy project permitting Zhang et al. (2024). By leveraging human expertise and domain knowledge, curated question sets can encompass a broader spectrum of linguistic variations, domain-specific considerations, and nuanced semantics Ribeiro et al. (2020), providing a more comprehensive evaluation of RAG’s performance across diverse scenarios and applications Thakur et al. (2021). Combining automated generation with human curation for benchmarking RAG offers a synergistic approach to ensure both efficiency and quality in question sets. This hybrid approach leverages the strengths of both automated and human-driven processes, that provide efficient and robust evaluation metrics for RAG’s performance.
@@
In this work, we present a hybrid workflow to benchmark RAG s, which combines rapid question generation through automated methods, augmented with properly designed human prompts to generate diverse set of questions. Our proposed benchmarking framework is used to generate questions from EIS and other research documents related to environmental impact of wind energy projects. The extensive question-answer dataset serve as a tool to evaluate the performance of RAG-based LLM s, which are designed to answer queries related to these extensive and comprehensive documents. Given the vast amount of information contained in these documents, manually reviewing them is impractical, making RAG-based LLM s essential for generating accurate responses to specific queries. Our benchmarking framework assesses the effectiveness of these models in accurately retrieving and responding to queries, ensuring that they can reliably process and provide relevant information from the documents.
Contributions  The paper introduces a novel benchmark dataset for question-answering (QA) task in a specific domain and also proposes a generic framework to evaluate the RAG-based LLM responses to different entries in the benchmark. This framework is designed to be adaptable across various domains, with a specific focus on documents related to wind energy project permitting in this study. The contributions of this research are as follows:
Novel domain-specific benchmark.
We present WeQA, [^1] This benchmark will be made publicly available. the first comprehensive benchmark QA dataset specifically designed for the wind energy domain, addressing the gap in specialized evaluation datasets for wind energy project permitting.
Domain-agnostic framework.
Our proposed benchmark creation and LLM evaluation framework is domain-agnostic and can be tailored for any desired niche domain, enabling researchers to adapt the methodology for various specialized fields beyond wind energy.
Hybrid question generation.
We introduce a hybrid method that automatically generates diverse question types with varying complexity levels, producing both objective and subjective responses across different document sections to comprehensively evaluate LLM performance.
Scalable evaluation methodology.
We utilize established scoring frameworks like RAGAS Es et al. (2023) and incorporate multiple LLM s as judges, ensuring scalability, reproducibility, and comprehensive performance assessment of RAG-based systems.
@@ 2 Related Works
General QA benchmarks.  These benchmarks have established foundational evaluation frameworks for reading comprehension and knowledge retrieval tasks. Notable general QA benchmarks include reading comprehension datasets such as the Stanford Question Answering Dataset (SQuAD) Rajpurkar et al. (2016) and MCTest Richardson et al. (2013), reasoning-focused benchmarks like the AI2 Reasoning Challenge (ARC) Clark et al. (2018), and comprehensive evaluation suites such as GLUE Wang et al. (2018) and Big Bench Srivastava et al. (2022). Additional benchmarks targeting open-domain knowledge include CommonsenseQA Talmor et al. (2018), TriviaQA Joshi et al. (2017), Search QA Dunn et al. (2017), and NewsQA Trischler et al. (2016).
Domain-specific QA benchmarks.  Recognizing the limitations of general benchmarks for specialized applications, researchers have developed domain-specific evaluation frameworks that capture the unique linguistic patterns, technical terminology, and reasoning requirements of particular fields. While scientific benchmarks such as MMLU Hendrycks et al. (2020), SciBench Wang et al. (2023), SciQ Welbl et al. (2017), SciRepEval Singh et al. (2022), SciQA Auer et al. (2023), and QASA Lee et al. (2023) are used for multi-disciplinary scientific QA evaluations, field-specific benchmarks include TheoremQA Chen et al. (2023c) for mathematics, emrQA Pampari et al. (2018) for medicine, BioRead Pappas et al. (2018) and BioMRC Pappas et al. (2020) for biology, LawBench Chen et al. (2023b) for legal, and NuclearQA Acharya et al. (2023) for nuclear domains.
For environmental assessment specifically, benchmarks such as EnviroExam Huang et al. (2024) for environmental science QA and NEPAQuAD Phan et al. (2023) for Environmental Impact Statement (EIS) documents have emerged. However, to our knowledge, no benchmarks exist specifically for wind energy project permitting, making the proposed WeQA benchmark the first comprehensive benchmarking effort in this critical domain.
@@ 3 Dataset Creation
We constructed a data extraction and curation pipeline to extract text, image, and table information from wind energy-related documents as depicted in the ‘data curation pipeline’ in Figure 1. Utilizing large language model (LLM) based methods such as the Unstructured.io tool Raymond (2023), we efficiently extracted information and converted it into JSON elements. To ensure data quality, we implemented a filtering step to remove images without meaningful content, such as decorative elements or blank spaces. These filtered JSON elements were then organized into a schema, creating a page-wise assortment of text, table, and image elements. This structured format ensures that the extracted data is easily accessible and can be accurately referenced during model training and evaluation.
@@ 4 Methodology
Figure 2: Different types of questions generated from the “introduction” section of a report Invenergy (2014) generated by the hybrid context approach. The section from the original document is first summarized and the question-answer pairs are generated from the summarized text chunk.
We present two complementary approaches for hybrid question generation to support comprehensive LLM benchmarking. The Hybrid Prompt Approach employs engineered prompts to generate high-quality, curated questions, while the Hybrid Context Approach leverages text summarization to create questions that require broader contextual understanding. The detailed prompts used for question generation across both approaches are provided in the Appendix.
Hybrid Prompt Approach. We utilize GPT-4 to automatically generate questions from given text chunks through carefully designed prompts tailored to each question type. To enhance question quality, we implement a manual curation process where domain experts identify exemplary questions that effectively assess LLM capabilities for benchmarking purposes. This curation is performed systematically across all question types, ensuring that each category incorporates appropriate grammatical structures and complexity levels. These curated questions subsequently serve as few-shot examples to guide the automatic question generation framework, improving the overall quality and consistency of generated questions.
Hybrid Context Approach. The initial approach primarily generates questions at the sentence level by substituting subjects or objects with interrogative words, which proves adequate for ‘closed’, ‘open’, and ‘recall’ type questions where answers can be directly extracted from the text. However, ‘process’, ‘evaluation’, and ‘comparison’ questions require deeper inferential reasoning across larger text segments. To address this limitation, we first employ GPT-4 to summarize extensive text chunks (typically exceeding 15 sentences) into concise summaries containing 5-8 sentences. We then generate questions from these summarized chunks using the hybrid prompt methodology combined with curated sample questions, ensuring that the resulting questions necessitate comprehensive understanding and synthesis of broader contextual information.
Questions from tables.  An essential component of benchmarking RAG-based LLM s within research articles and reports involves evaluating their capability to retrieve and interpret tabular information. Tables represent critical content elements within research documents, frequently containing comprehensive summaries and key quantitative data that encapsulate the essence of entire sections or studies. To address this requirement, we extract tabular data in HTML format and systematically organize it within our JSON schema framework. This HTML-formatted tabular data is subsequently incorporated into our prompt engineering pipeline to generate targeted question-answer pairs that specifically assess the model’s proficiency in understanding and reasoning over structured tabular information.
Figure 2 illustrates the diverse question-answer pairs generated from the introduction section of a document Invenergy (2014) using our proposed methodology. We demonstrate the Hybrid Context approach where the section content is first summarized into a concise form, and subsequently, targeted QA pairs are generated from this summarized context to ensure comprehensive coverage of key concepts. Table 1 presents the statistical distribution of different question types within the WeQA benchmark, providing insights into the composition and balance of our evaluation dataset.
Table 1: Question types in the WeQA benchmark
Type #Questions % Questions
Closed 382 18%
Comparison 393 19%
Evaluation 273 13%
Rhetorical 324 16%
Process 172 8%
Recall 258 12%
Open 270 13%
5 Results and Discussion
Experimental setup. We conduct a comprehensive evaluation of three state-of-the-art LLMs—GPT-4, Gemini, and Claude—on our WeQA benchmark within a RAG framework Knowledge extraction is performed from wind energy documents to create vector embeddings as shown in the data-curation pipeline in Figure 1, which are subsequently stored in a vector database to enable retrieval-augmented generation capabilities. We employ the $\mathsf{RAGAS}$ evaluation framework, leveraging judge LLMs to provide systematic assessment of model performance across multiple dimensions. The evaluation encompasses key metrics including answer correctness, context precision, and context recall, offering comprehensive insights into each model’s proficiency in both retrieving relevant information and generating accurate responses from the provided context. For the judge LLM component, we utilize both GPT-4 and Gemini-1.5Pro to ensure robust and unbiased evaluation of the assessed models’ performance. Figure 3 presents the answer correctness score, while the context precision and context recall depicted in Table 3 (added in Appendix) show the ability of the models to retrieve the context accurately.
Observation 1
Observation 2
There is an alignment in evaluations made by the two judge LLM s used within the $\mathsf{RAGAS}$ framework, particularly visible for ‘closed’ type questions.
This similarity arises because the answers to these questions are objective (‘yes’ or ‘no’), leading to equivalent correctness evaluations by both models. Although there are some mismatches in the evaluations made by the two judge LLM s, the number of these discrepancies is insignificant compared to the number of matching evaluations.
Figure 3: Answer correctness scores computed using the RAGAS scoring framework with GPT-4 and Gemini-1.5Pro as judge models for response generated by all three models used.
Figure 4 displays the confusion matrix illustrating the evaluations made by the two judge LLM s (GPT-4 and Gemini-1.5Pro) on the responses provided by the RAG-based Claude and GPT-4 models to the benchmarking questions. In this context, a true positive occurs when the judge LLM correctly identifies the model response as matching the ground truth. Conversely, a false positive arises when the judge LLM incorrectly states that the model response matches the ground truth, while it does not. This matrix helps visualize the accuracy and reliability of the evaluations conducted by the LLM s, when used within the $\mathsf{RAGAS}$ framework. We note that majority of evaluations made by either judge LLM matches the actual evaluation which indicates that both of them are reliable.
Observation 3
From this observation, we conclude that RAG-based models generate more accurate subjective responses to ‘open’ questions than objective (‘yes’ or ‘no’) responses for ‘closed’ questions. This phenomenon may stem from the inherent design of LLM s, which are optimized for generating extensive text sequences and may struggle with the precision required for definitive binary responses. This suggests that these models perform better when tasked with generating detailed, context-rich answers rather than simple, binary ones, highlighting their strength in handling nuanced and complex queries.
Figure 4: Confusion Matrix for evaluations by judge LLM s on responses from Claude (top) and GPT-4 (bottom) models
Observation 4
Observation 5
Observation 6
Figure 5 shows the scores computed by the judge LLM s for the responses generated by the three RAG-based LLM s. The Gemini-1.5Pro judge tends to rate high scores even when the LLM s refuse to answer. An example is listed below:
Evaluation comparison of LLM generated answer Question: In the HTML table that estimates the annual number of bird collisions at different percentages of avoidance, what is the estimated number of collisions per year at $98.0\%$ avoidance? Expected answer: The estimated number of collisions per year at $98.0\%$ avoidance is $152$. Gemini generated answer: This question cannot be answered from the given source. While the text discusses bird collisions with wind turbines, it does not provide any data on estimated collision numbers at different avoidance rates. GPT-4 score: 0.194 Gemini-1.5Pro score 0.813
Conversely, in instances where the LLMs generate correct answers, Gemini-1.5Pro has been observed to evaluate them as incorrect; such as:
Evaluation comparison of LLM generated answer Question: Who is the GIS Technician in the ‘STUDY PARTICIPANTS’ table? Expected answer: JR Boehrs Gemini generated answer: Saif Nomani JR Boehrs was the GIS Technician. GPT-4 score: 0.703 Gemini-1.5Pro score: 0.200
Figure 5: Answer correctness (top left), answer similarity (top right), context precision (bottom left) and recall (bottom right) scores across different judge and generator models.
6 Conclusion
In conclusion, this paper presents a versatile framework for evaluating the performance of RAG-based LLM s across various question types and document sections. We showcase this by introducing a hybrid, automated question-generation method that ensures comprehensive coverage of both objective and subjective queries, and implement this for the use case of wind energy related document and present the WeQA benchmark, which is a first of its kind benchmark in wind energy domain. However, the usefulness of our work goes beyond this niche domain as our approach is domain-agnostic, meaning it can be used for creating benchmark for any domain. Additionally, our use of the RAGAS scoring framework allows for a thorough evaluation of model performance, offering a holistic assessment of LLM capabilities, while also having the advantage of being easy for other researchers to adapt this approach for their own work.
7 Limitations
A limitation of the proposed framework is that the automatic method of generating questions often produces queries that are too specific to the document from which they were derived. When these questions are posed to an LLM with a large document corpus, the model may struggle to respond accurately, necessitating the filtering of ambiguous questions to ensure relevance and clarity. Additionally, the $\mathsf{RAGAS}$ scoring framework, which relies on LLM s as judges, introduces uncertainty in performance metrics, as different judge LLM s may score responses differently. While comparisons can be made for questions with objective responses, evaluating and comparing subjective responses across different LLM s remains challenging and less consistent. Another limitation of this study is the absence of comprehensive ablation studies, including comparisons between RAG-enabled and non-RAG configurations, which would provide deeper insights into the specific contributions of retrieval mechanisms to model performance.
8 Ethical Considerations
Appendix A Prompts used to generate QA pairs using Hybrid Prompt Approach
In this section, we detail the various prompts used to create the different types of questions in the WeQA benchmark dataset. First, we show the prompt to generate questions from a given text chunk. We use curly braces to denote placeholders for the different inputs to the prompt.